OpenStudy (anonymous):

How to simplify 25^2/3

OpenStudy (anonymous):

25^2/3= 625/3= 208.3333333 reduced to 208.3 with 3 repeating

OpenStudy (anonymous):

Sorry if I wasn't clear, look at the drawing below.

OpenStudy (anonymous):



OpenStudy (anonymous):

oh so itll be